OpenStudy (anonymous):

Problem set 2, 5A-1: What does evaluate mean in terms of giving out the right answer? Using book of tables or calculator for direct answer? Or deriving the y from tan y = sqrt(3) with geometric proof?

OpenStudy (anonymous):

It means solving for for y. So yes, you derive y from tan y = sqrt(3), but you don't use a "geometric" proof. You look at your helpful Unit Circle and find where a sin/cos is equal to sqrt(3). At pi/3 you'll have a sine function equal to sqrt(3)/2 and a cosine function equal to 1/2. If you divide sine by cosine, you get sqrt(3). Let me know if you have any questions.

OpenStudy (anonymous):

Thanks. This was exactly what I was looking for, I knew what asin and acosine function values were based on unit circle, but was not sure if it was meant to prove also cos x = 1/2 -> x = pi/3, which proved to be more trickier than I thought. got it via: if tan x = sqrt(3), then hypotenuse is 2 and therefore cos x = 1/2.